Let x_1, x_2, \dots, x_{100} be real numbers. If
x_1^2 + 2x_2^2 + \dots + 100x_{100}^2 = 1,
then find the maximum value of x_1/1 + x_2/2 + \dots + x_{100}/100.

Answer by CPhill(2189)

Let $a_k = x_k \sqrt{k}$ for $k = 1, 2, \dots, 100$.
Then the given condition becomes
$$\sum_{k=1}^{100} x_k^2 k = \sum_{k=1}^{100} a_k^2 = 1$$
We want to find the maximum value of
$$S = \sum_{k=1}^{100} \frac{x_k}{k} = \sum_{k=1}^{100} \frac{a_k}{k\sqrt{k}}$$
By the Cauchy-Schwarz inequality, we have
$$\left( \sum_{k=1}^{100} \frac{a_k}{k\sqrt{k}} \right)^2 \le \left( \sum_{k=1}^{100} a_k^2 \right) \left( \sum_{k=1}^{100} \frac{1}{k^3} \right)$$
Since $\sum_{k=1}^{100} a_k^2 = 1$, we have
$$S^2 \le \sum_{k=1}^{100} \frac{1}{k^3}$$
$$S \le \sqrt{\sum_{k=1}^{100} \frac{1}{k^3}}$$
Thus, the maximum value of $\sum_{k=1}^{100} \frac{x_k}{k}$ is $\sqrt{\sum_{k=1}^{100} \frac{1}{k^3}}$.
To achieve equality in Cauchy-Schwarz inequality, we need
$$\frac{a_1}{1\sqrt{1}} = \frac{a_2}{2\sqrt{2}} = \dots = \frac{a_{100}}{100\sqrt{100}} = c$$
where $c$ is a constant.
Then $a_k = ck\sqrt{k}$, so $x_k \sqrt{k} = ck\sqrt{k}$, which means $x_k = ck$.
Substituting into the given condition, we have
$$\sum_{k=1}^{100} kx_k^2 = \sum_{k=1}^{100} k(ck)^2 = c^2 \sum_{k=1}^{100} k^3 = 1$$
$$c^2 = \frac{1}{\sum_{k=1}^{100} k^3}$$
$$c = \frac{1}{\sqrt{\sum_{k=1}^{100} k^3}}$$
Then $x_k = \frac{k}{\sqrt{\sum_{k=1}^{100} k^3}}$.
Therefore, the maximum value of $\sum_{k=1}^{100} \frac{x_k}{k}$ is $\sqrt{\sum_{k=1}^{100} \frac{1}{k^3}}$.
Final Answer: The final answer is $\boxed{\sqrt{\sum_{k=1}^{100} \frac{1}{k^3}}}$